# Modeling Saliency Dataset Bias

### v1

@@ 1 Introduction
To address this issue, we propose a saliency model that is able to account for dataset biases with less than 20 interpretable parameters. Adapting only the bias parameters to new data closes about 80% of the generalization gap and can be done on as little as 50 samples. This allows us to substantially improve the state-of-the-art on the MIT300[23], CAT2000[2] and COCO-Freeview[8] benchmarks in generalization, adaptation and full-training settings. Furthermore, analyzing the learned dataset specific parameters provides valuable insights into the variability of saliency across datasets. Our main contributions are as follows:
@@
Figure 1: (a) Model Architecture: An input image is rescaled into different resolutions, some defined in total image size in pixels, others in pixels per degree of visual angle. For each image, deep activations from CLIP and DINOv2 encoders are extracted and averaged across scales, from which a priority map is decoded which is then postprocessed with Blur, priority scaling and centerbias. (b) We combine five different large datasets to train and evaluate our models.
@@ 2 Related Works
The SALICON model [18] uses a two-scale backbone. Here, we extend and adapt this approach to use a substantially larger number of scales of absolute and relative size, and we don’t concatenate the features from multiple scales, but average them. EML-NET [21] uses multiple layers from multiple backbones, which is related to our approach of combining deep features from CLIP and DINOv2 backbones. The UNISAL model [12] trains a saliency model jointly for video and image saliency datasets and uses some domain specific modules to adapt to the different domains. Here we focus only on combining multiple image saliency datasets and analysing the differences within this domain. Furthermore, in UNISAL, most domain specific parameters are not interpretable (batch norm adaptation and linear weights for deep features). Finally, while both UNISAL and we learn domain or dataset specific parameters, our number of parameters is magnitudes smaller, enabling data-efficient adaptation to new datasets. SalFBNet [11] introduced feedback connections from deeper layers of the encoder network to early layers to allow adaptation of low-level features with high-level knowledge. In addition, they pretrained the model on a pseudo saliency dataset obtained by averaging the predictions of top performing saliency models on a large set of images. DeepGaze IIE [35] employed an ensembling strategy (see also [59]) to average predictions from multiple internal saliency models using different encoder backbones and showed that this results in high performance and good confidence calibration on new datasets. DeepGaze IIE currently represents the state of the art in all datatsets and benchmarks presented in this paper. Most saliency models use convolutional architectures which keeps the spatial information in a very direct way, more recently also transformer architectures have been introduced [39, 65]. Here, we combine convolutional and transformer based backbones.
@@ 3 Model
Our model architecture is visualized in Figure 1 and can be seen as an extension of the DeepGaze architecture [32, 35]. Firstly, an input image is rescaled into different resolutions. Secondly, for each resolution, deep activations from CLIP[45] and DINOv2[41] encoders are extracted. CLIP is very good at encoding global information about a scene [67, 24, 37], while DINOv2 is designed to encode very precise local information and together they should allow to reason about objects within context. Thirdly, the activations from all scales are scaled into a common resolution and combined into a weighted average. Fourthly, the resulting feature maps are decoded into a single spatial priority map with a readout network [32], consisting of five layers of 1x1 convolutions. The small number of parameters makes the readout network trainable on datasets with only 1000 images. Lastly, the spatial priority map is scaled with a priority scaling factor, blurred with a Gaussian and combined with a precomputed center bias log density map which can be down weighted with a weight factor. The result is converted into a probability distribution over pixels using a softmax. For more details on the model architecture, see Appendix G.
@@
Nearly all parameters of our model are encompassed by the readout network with a total of 26,460 parameters which are trained jointly across datasets. The remaining parameters are what we call dataset bias parameters: they are dataset specific and control interpretable mechanisms. The dataset bias parameters are: the multiscale averaging weights which constitute 10 dataset bias parameters that model the relationship between how large the object is both relative to the image but also in terms of how large the object is perceived visually [44]; the priority scaling models how much more salient, e.g. a face is compared to a house, and we assume it will vary depending on the experimental conditions, e.g., the engagement of the observers; the blur size specifies the size of the blur kernel (specified in dva) as we assume that datasets might differ in how much fixations are spread out around objects; the center bias constitutes 2-5 dataset bias parameters depending on how exactly the center bias is modeled (see Appendix D for more details). It is the spatial prior of the model and encodes a tendency to fixate more towards the center of the image [54]. Lastly, the center bias weight accounts for the fact that a part of the empirical centerbias from the data, which we use in our models as spatial prior, is explained from image content [13] and hence already predicted by the readout network. Hence, we allow our models to downweight the center bias and make this weight dataset dependent as it depends on the image selection in the datasets. In total, this amounts to fewer than 20 dataset bias parameters with interpretable values. We refer to model variants where the bias parameters are identical across datasets as naive, as opposed to the full, bias-aware model.
@@ 4 Experiment Setup
We use five different datasets in our study: MIT1003 [23], CAT2000 [2], COCO Freeview [8, 65], DAEMONS [53] and FIGRIM [4]. Where included, we discard the initial forced fixation. Unlike all other datasets, FIGRIM technically is not freeviewing data since subjects are doing an image reidentification task where they have to identify repeated image presentations. However, in free-viewing experiments, subjects are commonly told that they will later have to re-identify images to keep engagement up. Hence, we assume that in FIGRIM, the eye movement data for the first presentation of each image to a subject is essentially freeviewing data. We exclude eye movement data from repeated image presentations. We noticed that the CAT2000 dataset contains an artifact, which we filtered out (see Appendix K). For more details on the used datasets including validation splits see Appendix H.
@@
Since our model computes a 2d fixation probability distribution, we can compute average log-likelihood for ground truth fixations by taking the log-probability of the pixel for each fixation and averaging values. Log-likelihood has been shown to be a very powerful loss function for saliency models that generalizes very well to all commonly used saliency metrics [31, 33, 28]. All our models are first pretrained on the SALICON dataset [22] and then trained in the actual training setting. Overall, we have 4 main training setups: (1) we train one model individually on each dataset and evaluate it on all datasets; for each dataset we train (2) one bias-naive model on the four other datasets (leave-one-out setting) and evaluate the target dataset and (3) one bias-aware model which is evaluated with averaged or adapted dataset parameters; (4) lastly, we train the full model on all five datasets. For more details on the training see Appendix I.
@@
We include two baseline models. The centerbias model is a KDE which, for each image, uses the fixation locations from all other images in the dataset. The centerbias quantifies how well fixations can be predicted without taking image content into account.The gold standard model estimates inter-observer consistency and is implemented as suggested by [28], as a regularized KDE which is crossevaluated in a leave-one-subject-out-setting. In some figures, we specify the gold standard as a range with the upper limit being the mixture of all observer’s gold standard models. For more details on the baseline models, see Appendix J. For comparing to previous state-of-the-art we include DeepGaze IIE [35], UNISAL [12], SalFBNet [11] and EML-Net [21]. Each of the models is applied in the resolution which resulted in the highest performance on each the dataset.
@@ 5 Results
In Fig. 2 we show our main results: Compared to full training on all five datasets, we find that naively generalizing from one dataset to another one results in a substantial performance drop (inter-dataset gap) of more than 40% information gain. Opposed to what one might expect, this gap cannot be closed by simply training on more diverse data: When training naively on four other datasets, 58% of the inter-dataset-gap still remains: there is a substantial generalization gap to unseen datasets. By adapting just the dataset-specific parameters of our bias-aware model to the target dataset, we close 76% of the generalization gap and reach close to full performance. For all training setups we report averages over mean scores from all five datasets. In Sec. 5 we show performances per dataset and also evaluate AUC.
@@
We test how well we can adapt the dataset bias parameters to unseen datasets when using limited data in our leave-one-dataset-out setting. To this end, we adapt the dataset parameters on small random subsets of the target dataset and evaluate on the full validation split. Due to the small sample sizes, for the model centerbias, we use a mixture of a KDE, a centered Gaussian and a uniform component (for more details see Appendix D).
As shown in Fig. 4 we outperform the generalization case with as little as 5 to 10 images for finetuning. With 50 images we already perform close to the full adaptation score. This demonstrates that the dataset bias parameters can successfully be adapted using minimal data, which is an important consideration for real world applications. Since all our dataset bias parameters are interpretable, it may even be possible in some use cases to infer them without any data, yielding performance that surpasses merely averaging the biases of all training datasets.
@@
Opposed to the leave-one-out-setting before, in Fig. 3, we compare for each dataset the performance of our jointly trained model with dataset biases with two other training setups: the model is trained jointly on all datasets but without the dataset-specific bias parameters (left) and the model trained separately for each dataset (right). We see that the joint model without dataset-specific bias parameters performs worse on all datasets compared to models trained separately. This finding supports our hypothesis that datasets differ in ways that overshadow the benefits of better learned patterns due to more diverse training data. Adding dataset specific bias parameters to the jointly trained model compensates for this problem and results in performance comparable to individually trained models. The fact that the joint bias-aware model does not yet outperform individually trained models suggests that additional dataset biases, e.g. semantic biases, might be at play which our model does not capture yet.
@@
1.3239 0.8982 - 0.4730 0.8840 0.6930 0.8673 0.8829 -
We test our model on the three datasets of the MIT/Tuebingen Saliency Benchmark [29]: MIT300 (the test set for MIT1003), CAT2000 and COCO-Freeview in three different setups: generalization and adaptation of the bias-aware model trained on the respective four other training datasets, and full joint training. For all models and datasets, we set a new state-of-the-art. In particular, with full joint training increasing the main ranking metric AUC by at least 1.1%–1.5% in all datasets, with adaptation a close second, emphasizing the power of our modeling approach (Sec. 5, also Appendix A for all metrics).
@@ 6 Analyses and Ablations
Figure 5: Contribution of different biases to closing the generalization gap. Performances are averages across the five datasets. Error bars [9, 40] are quite large because contributions differ across different datasets (Appendix, Fig. 13)
In order to understand which of the different dataset biases implemented in our model are most relevant for closing the generalization gap, we conduct an ablation study in the leave-one-dataset-out setting: We start evaluating each target-dataset in the generalization setting with averaged bias parameters. We then adapt more and more bias parameters on the target dataset to see how performance increases. Results averaged across all five datasets are shown in Fig. 5. We see that centerbias and multiscale weights contribute equally and account for a large part of the performance gain. Priority scaling adds a bit more, the other effects are barely noticable in the dataset average. However, if we compare the performances separate for each dataset (Appendix Fig. 13), we see that which bias matters how much varies from dataset to dataset. MIT1003 and CAT2000 gain from adapting the centerbias weight, and DAEMONS and FIGRIM gain from adapting the blur size, showing that each bias effect is useful for some datasets.
@@
In Fig. 6, we analyse example predictions from cases where our model outperforms the previous state-of-the-art DeepGaze IIE most (success cases), or where it misses most performance compared to inter-observer consistency (failure cases). We find that our model excels at predicting fixations at hidden faces, objects like instruments and toy animals and gets less distracted by some patterns. This shows that our model not only quantitatively, but also qualitatively improves over DeepGaze IIE. In terms of failure cases, we find that the model underestimates the saliency of low level pattern and abstract drawings, overestimates the salience of background objects and misses fixations on occluded objects. Fixing these issues would both require better high-level understanding of scenes as well as a better understanding of the interplay between high-level and low-level visual features and demonstrates that spatial saliency is still not solved.
@@
In Fig. 7a we compare the performance of different models which step by step transform the previous SOTA model DeepGaze IIE into our new model to quantify the contribution of the different architectural elements. We see that each of the three main architectural changes (replacing the ensemble over multiple backones with a combined CLIP+DINOv2 backbone, adding a multiscale feature extraction and finally making the bias parameters dataset dependent) contributes to the increased performance of our new model.
@@
To make sure that our claims about inter-dataset gap, generalization gap and closing thereof with few dataset biases do not depend on our specific backbone (CLIP+DINOv2), we run the same experiments with a variety of different backbones. The backbones are choosen to include the backbones from DeepGaze I [30], DeepGaze II [32] and some of the backbones used in DeepGaze IIE [35]. The results are shown in Fig. 7b and confirm that our results hold across backbones.
@@
In Fig. 7c we show that generalization from a naive-model, generalization from the bias-aware model and generalization via ensembling of per-dataset trained models perform roughly on par and far worse than the bias adaptation. Which generalization strategy works best differes from dataset to dataset.
@@
In Figure 8 we show how some of the dataset specific parameters differ across datasets for the model jointly trained across all five datasets. In Figure 8a we see that all datasets require both absolute and relative scales but the specific differ substantially across datasets. DAEMONS requires very high-resolution scales, FIGRIM profits from low-resolution relative scales and most other datasets are roughly in the middle. In Figure 8b we show the learned priority scaling parameters across dataset and as a function of presentation time. Again the learned values differ substantially. There is a clear dependency on presentation time, as, e.g., also suggested by Schütt et al. [50], but also variance beyond that (as visible for CAT2000 vs COCO-Freeview). We also find that the Gaussian blur differs substantially across datasets with DAEMONS requiring much smaller blur size, and that the centerbias weights differ across datasets (Appendix Fig. 17).
@@
In the Appendix we include additional analyses: We conduct a sensitivity analysis for the bias parameters (Appendix B), we test generalization and adaptation on the Toronto, Kienzle and SALICON datasets (Appendix C), we explore differences between models trained across datasets or individually (Appendix F), and we compare different strategies for modeling the center bias (Appendix D)

### v2

@@ 1 Introduction
To address this issue, we propose a saliency model that is able to account for dataset biases with less than 20 interpretable parameters. Adapting only the bias parameters to new data closes about 75% of the generalization gap and can be done on as little as 50 samples. This allows us to substantially improve the state-of-the-art on the MIT300 [27], CAT2000 [4] and COCO-Freeview [10] benchmarks in generalization, adaptation and full-training settings. Furthermore, analyzing the learned dataset specific parameters provides valuable insights into the variability of saliency across datasets. Our main contributions are as follows:
@@
Figure 1: Model Architecture: An input image is rescaled into different resolutions, some defined in total image size in pixels, others in pixels per degree of visual angle. For each image, deep activations from CLIP and DINOv2 encoders are extracted and averaged across scales, from which a priority map is decoded which is then postprocessed with Blur, priority scaling and centerbias. See Appendix Figure 9 for a larger version.
@@ 2 Related Works
The SALICON model [22] uses a two-scale backbone. Here, we extend and adapt this approach to use a substantially larger number of scales of absolute and relative size, and we don’t concatenate the features from multiple scales, but average them. EML-NET [25] uses multiple layers from multiple backbones, which is related to our approach of combining deep features from CLIP and DINOv2 backbones. The UNISAL model [15] trains a saliency model jointly for video and image saliency datasets and uses some domain specific modules to adapt to the different domains. Here we focus only on combining multiple image saliency datasets and analysing the differences within this domain. Furthermore, in UNISAL, most domain specific parameters are not interpretable (batch norm adaptation and linear weights for deep features). Finally, while both UNISAL and we learn domain or dataset specific parameters, our number of parameters is magnitudes smaller, enabling data-efficient adaptation to new datasets. SalFBNet [13] introduced feedback connections from deeper layers of the encoder network to early layers to allow adaptation of low-level features with high-level knowledge. In addition, they pretrained the model on a pseudo saliency dataset obtained by averaging the predictions of top performing saliency models on a large set of images. DeepGaze IIE [39] employed an ensembling strategy (see also [65]) to average predictions from multiple internal saliency models using different encoder backbones and showed that this results in high performance and good confidence calibration on new datasets. DeepGaze IIE currently represents the state of the art in all datatsets and benchmarks presented in this paper. Most saliency models use convolutional architectures which keeps the spatial information in a very direct way, more recently also transformer architectures have been introduced [43, 72, 14]. Here, we combine convolutional and transformer based backbones. Recently, researchers tried to improve saliency prediction by adding time [2], Augmentations [3], explicit global semantic interactions [71] or self-knowledge distillation [60, 61]
@@ 3 Model
Our model architecture is visualized in Figure 1 and can be seen as a variant of the DeepGaze architecture [36, 39] that has been extended with more modern backbones, a multiscale architecture and where some parameters have been made dataset-dependent. Firstly, an input image is rescaled into different resolutions. Secondly, for each resolution, deep activations from pretrained backbones are extracted. Thirdly, the activations from all scales are scaled into a common resolution and combined into a weighted average. Fourthly, the resulting feature maps are decoded into a single spatial priority map with a readout network [36], consisting of five layers of 1x1 convolutions. The small number of parameters makes the readout network trainable on datasets with only 1000 images. Lastly, the spatial priority map is multiplied with a priority scaling factor, blurred with a Gaussian and combined with a precomputed center bias log density map which can be down weighted with a weight factor. The result is converted into a probability distribution over pixels using a softmax. In the following we detail our extensions to previous DeepGaze architectures. For a precise mathematical formulation and more details on the model architecture see Appendix H.
@@
Pretrained Backbones
For each resolution we extract deep activations from two different pretrained backbones which are then concatenated: CLIP [49] is very good at encoding global information about a scene [74, 28, 41], while DINOv2 [45] is designed to encode very precise local information. Together they should allow to reason about objects within context.
@@
Nearly all parameters of our model are encompassed by the readout network with a total of 26,460 parameters which are trained jointly across datasets. The remaining parameters are what we call dataset bias parameters: they are dataset specific and control interpretable mechanisms. The dataset bias parameters are: the multiscale averaging weights which constitute 10 dataset bias parameters that model the relationship between how large the object is both relative to the image but also in terms of how large the object is perceived visually [48]; the priority scaling models how much more salient, e.g. a face is compared to a house, and we assume it will vary depending on the experimental conditions, e.g., the engagement of the observers; the blur size specifies the size of the blur kernel (specified in dva) as we assume that datasets might differ in how much fixations are spread out around objects; the center bias constitutes 2-5 dataset bias parameters depending on how exactly the center bias is modeled (see Appendix E for more details). It is the spatial prior of the model and encodes a tendency to fixate more towards the center of the image [58]. Lastly, the center bias weight accounts for the fact that a part of the empirical centerbias from the data, which we use in our models as spatial prior, is explained from image content [16] and hence already predicted by the readout network. Hence, we allow our models to downweight the center bias and make this weight dataset dependent as it depends on the image selection in the datasets. In total, this amounts to fewer than 20 dataset bias parameters with interpretable values. We refer to model variants where the bias parameters are identical across datasets as naive, as opposed to the full, bias-aware model.
@@ 4 Experiment Setup
We use five different datasets in our study: MIT1003 [27], CAT2000 [4], COCO Freeview [10, 72], DAEMONS [57] and FIGRIM [6]. Where included, we discard the initial forced fixation. Unlike all other datasets, FIGRIM technically is not freeviewing data since subjects are doing an image reidentification task where they have to identify repeated image presentations. However, in free-viewing experiments, subjects are commonly told that they will later have to re-identify images to keep engagement up. Hence, we assume that in FIGRIM, the eye movement data for the first presentation of each image to a subject is essentially freeviewing data. We exclude eye movement data from repeated image presentations. We noticed that the CAT2000 dataset contains an artifact, which we filtered out (see Appendix L). For more details on the used datasets including validation splits see Appendix I.
@@
Since our model computes a 2d fixation probability distribution, we can compute average log-likelihood for ground truth fixations by taking the log-probability of the pixel for each fixation and averaging values. Log-likelihood has been shown to be a very powerful loss function for saliency models that generalizes very well to all commonly used saliency metrics [35, 37, 32]. All our models are first pretrained on the SALICON dataset [26] and then trained in the actual training setting. Overall, we have 4 main training setups: (1) we train one model individually on each dataset and evaluate it on all datasets; for each dataset we train (2) one bias-naive model on the four other datasets (leave-one-out setting) and evaluate the target dataset and (3) one bias-aware model which is evaluated with averaged or adapted dataset parameters; (4) lastly, we train the full model on all five datasets. For more details on the training see Appendix J.
@@
We include two baseline models. The centerbias model is a KDE which, for each image, uses the fixation locations from all other images in the dataset. The centerbias quantifies how well fixations can be predicted without taking image content into account.The gold standard model estimates inter-observer consistency and is implemented as suggested by [32], as a regularized KDE which is crossevaluated in a leave-one-subject-out-setting. In some figures, we specify the gold standard as a range with the upper limit being the mixture of all observer’s gold standard models. For more details on the baseline models, see Appendix K. For comparing to previous state-of-the-art we include DeepGaze IIE [39], UNISAL [15], SalFBNet [13] and EML-Net [25]. Each of the models is applied in the resolution which resulted in the highest performance on each the dataset.
@@ 5 Results
Table 1: Performance of our model and previous state-of-the-art models. Best performance in is indicated in bold, second best is underlined. “generalization” refers to training on the respective four other datasets and evaluation with average dataset biases, “adaptation” refers to training on the respective four other datasets and evaluation after adapting the dataset bias parameters to the target dataset. Models are sorted by average AUC. See App. B for an extended version with more comparison models.
@@
In Fig.˜2 we show our main results: Compared to full training on all five datasets, we find that naively generalizing from one dataset to another one results in a substantial performance drop (inter-dataset gap) of more than 40% information gain. Opposed to what one might expect, this gap cannot be closed by simply training on more diverse data: When training naively on four other datasets, 58% of the inter-dataset-gap still remains: there is a substantial generalization gap to unseen datasets. By adapting just the dataset-specific parameters of our bias-aware model to the target dataset, we close 76% of the generalization gap and reach close to full performance. For all training setups we report averages over mean scores from all five datasets. In Sec.˜5 we show performances per dataset and also evaluate AUC. In App. B we also compare with older saliency models, which on CAT2000 outperform some DNN based models.
@@
We test how well we can adapt the dataset bias parameters to unseen datasets when using limited data in our leave-one-dataset-out setting. To this end, we adapt the dataset parameters on small random subsets of the target dataset and evaluate on the full validation split. Due to the small sample sizes, for the model centerbias, we use a mixture of a KDE, a centered Gaussian and a uniform component (for more details see Appendix E).
As shown in Fig.˜4 we outperform the generalization case with as little as 5 to 10 images for finetuning. With 50 images we already perform close to the full adaptation score. This demonstrates that the dataset bias parameters can successfully be adapted using minimal data, which is an important consideration for real world applications. Since all our dataset bias parameters are interpretable, it may even be possible in some use cases to infer them without any data, yielding performance that surpasses merely averaging the biases of all training datasets.
@@
Opposed to the leave-one-out-setting before, in Fig.˜3, we compare for each dataset the performance of our jointly trained model with dataset biases with two other training setups: the model is trained jointly on all datasets but without the dataset-specific bias parameters (left) and the model trained separately for each dataset (right). We see that the joint model without dataset-specific bias parameters performs worse on all datasets compared to models trained separately. This finding supports our hypothesis that datasets differ in ways that overshadow the benefits of better learned patterns due to more diverse training data. Adding dataset specific bias parameters to the jointly trained model compensates for this problem and results in performance comparable to individually trained models. The fact that the joint bias-aware model does not yet outperform individually trained models suggests that additional dataset biases, e.g. semantic biases, might be at play which our model does not capture yet.
@@
MIT300 CAT2000 COCO-Freeview
Model IG AUC sAUC IG AUC sAUC IG AUC sAUC
TempSAL - 0.863 0.748 - 0.844 0.638 - 0.857 0.708
DeepGaze II 0.951 0.876 0.784 0.084 0.864 0.650 0.664 0.870 0.740
EML-NET - 0.876 0.747 - 0.831 0.585 - 0.845 0.707
SalFBNet 0.819 0.877 0.786 - 0.855 0.633 - 0.872 0.710
UNISAL 0.951 0.877 0.784 0.032 0.860 0.668 0.749 0.877 0.758
Clueify - 0.881 0.765 - - - - - -
DeepGaze IIE 1.071 0.883 0.794 0.189 0.869 0.668 0.860 0.882 0.767
Ours (generalized) 1.198 0.893 0.814 0.203 0.871 0.689 0.947 0.890 0.786
Ours (adapted) 1.236 0.894 0.815 0.433 0.881 0.690 1.011 0.893 0.788
Ours (full joint training) 1.246 0.894 0.816 0.493 0.885 0.700 1.073 0.897 0.795
Interobserver consistency 1.3239 0.8982 - 0.4730 0.8840 0.6930 0.8673 0.8829 -
We test our model on the three datasets of the MIT/Tuebingen Saliency Benchmark [33]: MIT300 (the test set for MIT1003), CAT2000 and COCO-Freeview in three different setups: generalization and adaptation of the bias-aware model trained on the respective four other training datasets, and full joint training. For all models and datasets, we set a new state-of-the-art. In particular, with full joint training increasing the main ranking metric AUC by at least 1.1%–1.5% in all datasets, with adaptation a close second, emphasizing the power of our modeling approach (Sec.˜5, also Appendix A for all metrics).
@@ 6 Analyses and Ablations
Figure 5: Contribution of different biases to closing the generalization gap. Performances are averages across the five datasets. Error bars [11, 44] are quite large because contributions differ across different datasets (Appendix, Fig.˜14)
In order to understand which of the different dataset biases implemented in our model are most relevant for closing the generalization gap, we conduct an ablation study in the leave-one-dataset-out setting: We start evaluating each target-dataset in the generalization setting with averaged bias parameters. We then adapt more and more bias parameters on the target dataset to see how performance increases. Results averaged across all five datasets are shown in Fig.˜5. We see that centerbias and multiscale weights contribute equally and account for a large part of the performance gain. Priority scaling adds a bit more, the other effects are barely noticable in the dataset average. However, if we compare the performances separate for each dataset (Appendix Fig.˜14), we see that which bias matters how much varies from dataset to dataset. MIT1003 and CAT2000 gain from adapting the centerbias weight, and DAEMONS and FIGRIM gain from adapting the blur size, showing that each bias effect is useful for some datasets.
@@
In Fig.˜6, we analyse example predictions from cases where our model outperforms the previous state-of-the-art DeepGaze IIE most (success cases), or where it misses most performance compared to inter-observer consistency (failure cases). We find that our model excels at predicting fixations at hidden faces, objects like instruments and toy animals and gets less distracted by some patterns. This shows that our model not only quantitatively, but also qualitatively improves over DeepGaze IIE. In terms of failure cases, we find that the model underestimates the saliency of low level pattern and abstract drawings, overestimates the salience of background objects and misses fixations on occluded objects. Fixing these issues would both require better high-level understanding of scenes as well as a better understanding of the interplay between high-level and low-level visual features and demonstrates that spatial saliency is still not solved.
@@
In Fig.˜7a we compare the performance of different models which step by step transform the previous SOTA model DeepGaze IIE into our new model to quantify the contribution of the different architectural elements. We see that each of the three main architectural changes (replacing the ensemble over multiple backones with a combined CLIP+DINOv2 backbone, adding a multiscale feature extraction and finally making the bias parameters dataset dependent) contributes to the increased performance of our new model.
@@
To make sure that our claims about inter-dataset gap, generalization gap and closing thereof with few dataset biases do not depend on our specific backbone (CLIP+DINOv2), we run the same experiments with a variety of different backbones. The backbones are choosen to include the backbones from DeepGaze I [34], DeepGaze II [36] and some of the backbones used in DeepGaze IIE [39]. The results are shown in Fig.˜7b and confirm that our results hold across backbones.
@@
In Fig.˜7c we show that generalization from a naive-model, generalization from the bias-aware model and generalization via ensembling of per-dataset trained models perform roughly on par and far worse than the bias adaptation. Which generalization strategy works best differes from dataset to dataset.
@@
In Figure 8 we show how some of the dataset specific parameters differ across datasets for the model jointly trained across all five datasets. In Figure 8a we see that all datasets require both absolute and relative scales but the specific differ substantially across datasets. DAEMONS requires very high-resolution scales, FIGRIM profits from low-resolution relative scales and most other datasets are roughly in the middle. In Figure 8b we show the learned priority scaling parameters across dataset and as a function of presentation time. Again the learned values differ substantially. There is a clear dependency on presentation time, as, e.g., also suggested by Schütt et al. [54], but also variance beyond that (as visible for CAT2000 vs COCO-Freeview). We also find that the Gaussian blur differs substantially across datasets with DAEMONS requiring much smaller blur size, and that the centerbias weights differ across datasets (Appendix Fig.˜18).
@@
In the Appendix we include additional analyses: We conduct a sensitivity analysis for the bias parameters (Appendix C), we test generalization and adaptation on the Toronto, Kienzle and SALICON datasets (Appendix D), we explore differences between models trained across datasets or individually (Appendix G), we assess relevance of different parts of the multiscale architecture in an ablation study (Appendix F) and we compare different strategies for modeling the center bias (Appendix E)